OpenStudy (anonymous):

Simplify. r^2s^5t^4/ r^2s^2t^5

OpenStudy (solomonzelman):

`Read what I say carefully ` v^(x) / v^(x+c) =1 / v^c v^(x+c) / v^(x) = v^c

OpenStudy (solomonzelman):

\[5=2+3~~~~~~~~~~~~~~~~and~~~~~~~~~~~~~~~5=4+1\]

OpenStudy (anonymous):

Oh ok. Thanks.

OpenStudy (jdoe0001):

\(\bf \cfrac{r^2s^5t^4}{r^2s^2t^5}\implies \cfrac{\cancel{ r^2s^{\color{brown}{ 2}} }s^{\color{brown}{ 3}}\cancel{ t^4 }}{\cancel{ r^2s^2 }\cancel{ t^{\color{brown}{ 4}} }t^{\color{brown}{ 1}}}\) as SolomonZelman pointed out